Mr. Paxton invested $2,000 more then his wife. the income from both investments at 5% was $250 in a year. How much did each invest?
Answer by funmath(2933)

Mr. Paxton invested $2,000 more then his wife. the income from both investments at 5% was $250 in a year. How much did each invest?
Interest=principle*rate*time (t=1, so we don't need to include it.)
Let the principle of the wife be: x
Then the principle of Mr. Paxon is: x+2000
The rate for both is 5/100=.05
The interest of the wife is: .05x
The interest of Mr. Paxon is: .05(x+2000)
If we add their interests we get 250
.05x+.05(x+2000)=250
.05x+.05x+100=250
.1x+100=250
.1x+100-100=250-100
.1x=150
.1x/.1=150/.1
x=1500
The wife invested: x=$1,500
Mr. Paxon invested: x+2000=$1,500+$2,000=$3,500
Sanity Check:
If you add the two prinples you get 1,500+3,500=5,000
Is 5% of 5000=250?
.05(5000)=250 yes. we're still sane.
Happy Calculating!!!!